# Targeted Error Correction in Knowledge Distillation: Small Language Models Surpass GPT

### v1

@@ 7 Results
We first evaluated the effectiveness of the "Analyze-Revise-Finetune" pipeline across various student LLMs. Table 2 presents the summarization performance for each model across different training conditions. Each model was assessed under four distinct scenarios: 1) without fine-tuning (out-of-box), 2) fine-tuned on the original training data (org), 3) fine-tuned on revision version 1 of the training data (r1), and 4) fine-tuned on revision version 2 of the training data (r2). To clearly illustrate the impact of revised training data, we include the score differences compared to the original dataset (org) for both r1 and r2 ($\Delta$ over org.).
@@
Table 2: Summarization performance across different training data sets and student LLMs. Mean of auto-evaluator rating (1-5 scale) is given. Superscript * marks performance better than teacher LLM (GPT-3.5). Best scores for each channel are bold-faced.
Bot Chat WebForm
Mean auto-rating $\Delta$ over org. Mean auto-rating $\Delta$ over org.
GPT3.5 (out-of-box) 4.05 4.29
Llama 3.1 8B (out-of-box) 2.34 3.32
Llama 3.1 8B (org) 4.14* 4.32*
Llama 3.1 8B (r1) 4.325* 0.185 4.405* 0.085
Llama 3.1 8B (r2) 4.315* 0.175 4.125 -0.195
Llama 3.1 70B (out-of-box) 4.28* 4.39*
Llama 3.1 70B (org) 3.96 4.365*
Llama 3.1 70B (r1) 4.315* 0.355 4.42* 0.055
Llama 3.1 70B (r2) 4.3* 0.34 4.22 -0.145
Lilium 2 7B (out-of-box) 1.145 1.975
Lilium 2 7B (org) 3.975 4.125
Lilium 2 7B (r1) 4.28* 0.305 4.295* 0.17
Lilium 2 7B (r2) 4.2* 0.225 4.185 0.06
Lilium 2 mix 7B (out-of-box) 1.015 1.47
Lilium 2 mix 7B (org) 3.825 4.125
Lilium 2 mix 7B (r1) 4.205* 0.38 4.11 -0.015
Lilium 2 mix 7B (r2) 4.16* 0.335 3.92 -0.205
Lilium 2 13B (out-of-box) 1.21 2.22
Lilium 2 13B (org) 3.755 3.995
Lilium 2 13B (r1) 4.075* 0.32 4.065 0.07
Lilium 2 13B (r2) 4.045* 0.29 3.8 -0.195
Lilium 2 mix 45B (out-of-box) 1.01 1.04
Lilium 2 mix 45B (org) 3.89 1.38
Lilium 2 mix 45B (r1) 4.225* 0.335 1.185 -0.195
Lilium 2 mix 45B (r2) 3.95 0.06 1.145 -0.235
Overall, the top-performing models were Llama 3.1 8B for the BotChat channel and Llama 3.1 70B for the WebForm channel, each fine-tuned using the r1 training set. Notably, these student models outperformed the GPT-3.5. This clearly demonstrates the effectiveness of the "Analyze-Revise-Finetune" method in enhancing smaller models to surpass larger, proprietary teacher models through targeted data revisions.
For the BotChat summarization task, all evaluated models exhibited performance improvements using both r1 and r2 revised training datasets, underscoring the benefit of our pipeline. Importantly, nearly all models fine-tuned on the revised datasets—except Lilium 2 mix 45B (r2)—exceeded the performance of GPT-3.5.This demonstrates that our "Analyze-Revise-Finetune" approach can effectively overcome the typical performance limitations associated with traditional knowledge distillation methods. In contrast, models fine-tuned solely on the original training set did not surpass GPT-3.5, with the single exception of Llama 3.1 8B.
In the WebForm summarization task, Llama 3.1 8B, Llama 3.1 70B, Lilium 2 7B, and Lilium 2 13B benefitted notably from the r1 training dataset, again surpassing GPT-3.5 and highlighting the effectiveness of fine-tuning through targeted error-correction for specific tasks. Interestingly, the two Lilium mixture-of-expert models (Lilium 2 mix 7B and Lilium 2 mix 45B) exhibited negative performance impacts when using the r1 dataset; further investigation into this phenomenon is left for future research.
However, for the WebForm summarization task, the r2 training dataset resulted in performance degradation for most student LLMs, except Lilium 2 7B. The reduced performance can be attributed to the quality of the revised summaries. As illustrated in Table 3, the revision success rate varied by error type. The revision addressing unn_content_webform_email_copyunn\_content\_webform\_email\_copy had a high success rate of 97%, generating the r1 dataset. In contrast, the revision targeting unn_content_redundantunn\_content\_redundant, creating the r2 dataset, had a success rate of only 61%. This lower revision quality also explains the smaller performance improvements seen in BotChat summaries when using the r2 dataset compared to the performance improvement using r1 dataset.
@@
Error type Bot Chat WebForm Overall
unn_content_requests_agent 94% - 94%
sentiment_inferred_frustrated 92% - 92%
unn_content_webform_email_copy - 97% 97%
unn_content_redundant 78% 61% 70%
Additionally, we explored how the sequence or order in which training data from different channels is provided affects summarization performance. Table 4 presents the summarization results for Llama 3.1 8B trained on the r1 dataset with various training sequences. Initially, we created two models, each trained exclusively on one channel: BotChat-only and WebForm-only. Then, these models were sequentially fine-tuned on data from the other channel, resulting in two additional models: one trained first on BotChat and then on WebForm (BotChat→WebForm), and another trained first on WebForm and then on BotChat (WebForm→BotChat). Finally, we trained a model using data from both channels simultaneously (BotChat+WebForm), which represents the default configuration for all other experiments discussed in this paper.
Table 4: Summarization performance of Llama 3.1 8B trained on r1 with different training sequences. Mean of auto-evaluator rating (1-5 scale) is given. Best scores for each channel are bold-faced.
Training sequence BotChat WebForm
Llama 3.1 8B (out-of-box) 2.34 3.32
BotChat only 4.265 3.33
WebForm only 2.85 4.215
BotChat→WebForm 2.57 3.405
WebForm→BotChat 2.77 3.41
BotChat+WebForm 4.325 4.405
For both channels, the best performance is achieved when training datasets from both channels are provided simultaneously (BotChat+WebForm). When only a single channel dataset is used for training (BotChat only, WebForm only), the performance for that channel remains similar but slightly lower compared to the BotChat+WebForm configuration. This suggests that despite channel-specific differences, the model benefits from cross-channel learning due to shared underlying principles and evaluation criteria for customer interaction summarization. Moreover, models trained exclusively on one channel (BotChat only, WebForm only) still show slight improvements in the summarization performance for the other channel when compared to the out-of-box Llama 3.1 8B performance, further confirming cross-channel knowledge transfer.
When training data is provided sequentially (BotChat→WebForm, WebForm→BotChat), mean auto-ratings are higher compared to out-of-box models for both channels, demonstrating that the models do gain some knowledge from sequential training. However, these ratings remain notably lower than those achieved through simultaneous training (BotChat+WebForm) or single-channel training (BotChat only, WebForm only). For the initially trained channel (BotChat in BotChat→WebForm, WebForm in WebForm→BotChat), there may have been performance degradation due to catastrophic forgetting [12, 13]. Conversely, the performance on the channel trained subsequently (WebForm in BotChat→WebForm, BotChat in WebForm→BotChat) also remains significantly lower than other training configurations, possibly due to reduced learning capacity following initial fine-tuning with LoRa.
@@ 8 Discussion
In this paper, we successfully fine-tuned a smaller student LLM to surpass the performance of a teacher LLM on customer service interaction summarization tasks. Through our proposed "Analyze-Revise-Finetune" approach, we demonstrated that Llama 3.1 8B could outperform GPT-3.5. This clearly illustrates the effectiveness of our pipeline for specialized applications, highlighting its potential for replacing large proprietary models with smaller open-source alternatives. Adopting smaller models can significantly mitigate cost and privacy concerns associated with larger, proprietary LLMs.
We established a detailed evaluation framework tailored specifically for production systems that summarize customer service interactions. This framework ensures a consistent and systematic evaluation of summary quality, allowing for precise identification of the most common errors requiring correction. Additionally, by separating the editor LLM from the teacher LLM and utilizing Llama 3.1 70B for training data revisions, our approach further optimizes both inference cost and data privacy, offering substantial advantages compared to approaches that directly employ large teacher models for revisions.
@@ 9 Conclusion
In this paper, we introduced and validated a fine-tuning pipeline incorporating targeted error-correction designed to enhance the quality of training data generated by large proprietary teacher models for smaller, open-source student models. By integrating detailed error analysis and targeted text revisions, our "Analyze-Revise-Finetune" approach achieves strong summarization performance while significantly reducing costs and addressing privacy concerns associated with proprietary systems.
The results suggest that the Analyze–Revise–Finetune pipeline can be successfully extended to a wide range of LLM-based applications, offering consistent improvements in cost efficiency, output quality, and data privacy.
@@ Acknowledgments
We’d like to thank Jillin Gerten, Vishal Khadake, Mehran Nadjarbashi Noghani, Swapnil Thorat, Ryan Wen, Zonghan Xu, Pooja Daryani, and Megha Korade for their support in this work.
@@ Appendix A Additional Table
Unnecessary information unn_content_redundant Summary includes unnecesary information, i.e., 1) redundant or repetitive information, 2) courtesy or plesantry, 3) bot’s response, 4) customer ranting, 5) lack of specific details in the original content, 6) requets for human agent, 7) transfer to next agent, 8) request for email copy of original webform
@@
unn_content_ebay_response_included

### v2

@@ 7 Results
We evaluated the effectiveness of the "Analyze-Revise-Finetune" pipeline across various student LLMs. Table 3 presents the summarization performance for each model across different training conditions. Each model was assessed under four distinct scenarios: 1) without fine-tuning (out-of-box), 2) fine-tuned on the original training data (org), 3) fine-tuned on revision version 1 of the training data (r1), and 4) fine-tuned on revision version 2 of the training data (r2). To clearly illustrate the impact of revised training data, we include the score differences compared to the original dataset (org) for out-of-box, r1 and r2 ($\Delta$ over org.).
@@
Table 3: Summarization performance across different training data sets and student LLMs. Mean of auto-evaluator rating (1-5 scale) is given. Superscript * marks performance better than teacher LLM (GPT-3.5). Best scores for each channel are bold-faced. r: mean of auto-rating, $\Delta$: difference over when trained on the original training dataset
Bot Chat Teammate Chat WebForm Email
r $\Delta$ r $\Delta$ r $\Delta$ r $\Delta$
GPT3.5 (out-of-box) $2.525$ $3.965$ $4.030$ $1.485$
Llama 3.1 8B (out-of-box) $1.630$ 0.850-0.850 $3.975^{*}$ $0.030$ $3.625$ 0.565-0.565 $1.185$ 0.400-0.400
Llama 3.1 8B (org) $2.480$ $3.945$ $4.190^{*}$ $1.585^{*}$
Llama 3.1 8B (r1) $3.435^{*}$ $0.955$ $4.370^{*}$ $0.425$ $\mathbf{4.325^{*}}$ $0.135$ $2.060^{*}$ $0.475$
Llama 3.1 8B (r2) $3.265^{*}$ $0.785$ $4.440^{*}$ $0.495$ $4.005$ 0.185-0.185 $\mathbf{2.062^{*}}$ $0.477$
Llama 3.2 3B (out-of-box) $1.195$ 0.595-0.595 $2.470$ 0.560-0.560 $2.425$ 0.580-0.580 $1.182$ 0.134-0.134
Llama 3.2 3B (org) $1.790$ $3.030$ $3.005$ $1.316$
Llama 3.2 3B (r1) $3.200^{*}$ $1.410$ $4.260^{*}$ $1.230$ $3.885$ $0.880$ $1.910^{*}$ $0.594$
Llama 3.2 3B (r2) $3.220^{*}$ $1.430$ $\mathbf{4.470^{*}}$ $1.440$ $3.875$ $0.870$ $1.721^{*}$ $0.405$
Llama 3.2 1B (out-of-box) $1.375$ 0.815-0.815 $1.400$ 2.075-2.075 $1.625$ 2.155-2.155 $1.000$ 0.721-0.721
Llama 3.2 1B (org) $2.190$ $3.475$ $3.780$ $1.336$
Llama 3.2 1B (r1) $3.030^{*}$ $0.840$ $4.040^{*}$ $0.565$ $3.905$ $0.125$ $1.600^{*}$ $0.264$
Llama 3.2 1B (r2) $3.165^{*}$ $0.975$ $4.035^{*}$ $0.560$ $3.720$ 0.060-0.060 $1.689^{*}$ $0.353$
Qwen3 4B (out-of-box) $2.345$ $0.250$ $4.080^{*}$ $0.395$ $3.975$ 0.035-0.035 $1.594^{*}$ $0.094$
Qwen3 4B (org) $2.095$ $3.685$ $4.010$ $1.500^{*}$
Qwen3 4B (r1) $3.000^{*}$ $0.905$ $4.150^{*}$ $0.465$ $4.115^{*}$ $0.105$ $1.633^{*}$ $0.133$
Qwen3 4B (r2) $3.065^{*}$ $0.970$ $4.350^{*}$ $0.665$ $3.835$ 0.175-0.175 $1.535^{*}$ $0.035$
Gemma3 1B (out-of-box) $1.265$ 0.700-0.700 $2.290$ 1.015-1.015 $1.320$ 2.460-2.460 $1.022$ 0.016-0.016
Gemma3 1B (org) $1.965$ $3.305$ $3.780$ $1.038$
Gemma3 1B (r1) $\mathbf{3.520^{*}}$ $1.555$ $3.235$ 0.070-0.070 $3.000$ 0.780-0.780 $1.027$ 0.011-0.011
Gemma3 1B (r2) $2.065$ $0.100$ $3.400$ $0.095$ $3.350$ 0.430-0.430 $1.130$ $0.092$
Gemma3 270M (out-of-box) $1.050$ 0.880-0.880 $1.170$ 1.415-1.415 $1.180$ 1.875-1.875 $1.014$ 0.008-0.008
Gemma3 270M (org) $1.930$ $2.585$ $3.055$ $1.022$
Gemma3 270M (r1) $2.120$ $0.190$ $2.645$ $0.060$ $2.675$ 0.380-0.380 $1.049$ $0.027$
Gemma3 270M (r2) $2.130$ $0.200$ $2.690$ $0.105$ $2.895$ 0.160-0.160 $1.022$ $0.000$
Phi3.5mini 4B (out-of-box) $1.610$ 0.305-0.305 $3.645$ 0.160-0.160 $3.080$ 0.890-0.890 $1.535^{*}$ $0.067$
Phi3.5mini 4B (org) $1.915$ $3.805$ $3.970$ $1.468$
Phi3.5mini 4B (r1) $2.285$ $0.370$ $4.435^{*}$ $0.630$ $3.960$ 0.010-0.010 $1.675^{*}$ $0.207$
Phi3.5mini 4B (r2) $2.290$ $0.375$ $4.305^{*}$ $0.500$ $3.775$ 0.195-0.195 $1.792^{*}$ $0.324$
Table 3 shows that revising the teacher-generated training set yields broad and substantial gains in student fine-tuning performance across channels. Relative to fine-tuning on the original dataset (“org”), training on the revised datasets improves BotChat performance for all student models (7/7): r1 increases the mean rating by +0.89 points on average (range +0.19 to +1.56), and r2 by +0.69 points (range +0.10 to +1.43). The effect is similarly consistent on TeammateChat, where r1 improves 6/7 models (mean +0.47) and r2 improves 7/7 (mean +0.55, reaching gains up to +1.44 for Llama 3.2 3B). Email summarization also benefits: r1 improves 6/7 models (mean +0.24) and r2 improves 6/7 (with 1/7 unchanged; mean +0.24). Overall, these results shows that targeted revisions to the distillation/fine-tuning data are a primary driver of downstream student quality, beyond simply performing fine-tuning on the original teacher outputs.
The revised datasets also enable student models to match or surpass the teacher more reliably across the channels. Under r1, both Llama 3.1 8B and Qwen3 4B outperform GPT-3.5 in all four channels simultaneously (e.g., Llama 3.1 8B r1: 3.435 vs 2.525 on BotChat, 4.370 vs 3.965 on TeammateChat, 4.325 vs 4.030 on WebForm, and 2.060 vs 1.485 on Email). Consistent with this, the best student score in each channel is achieved by a revised-data variant rather than “org” or out-of-box: Gemma3 1B (r1) is best on BotChat (3.520), Llama 3.2 3B (r2) is best on Teammate Chat (4.470), Llama 3.1 8B (r1) is best on WebForm (4.325), and Llama 3.1 8B (r2) is best on Email (2.062). This pattern indicates that dataset revision is not merely stabilizing performance, but is instrumental in reaching peak student capability.
Revisions are particularly valuable in cases where naive fine-tuning on “org” is counterproductive. For example, Qwen3 4B degrades from out-of-box to org on both BotChat (2.345 → 2.095) and Teammate Chat (4.080 → 3.685), but revised-data training recovers and substantially exceeds both baselines (Bot Chat 3.000–3.065, Teammate Chat 4.150–4.350). A similar recovery is visible for Phi3.5mini 4B on Email (org 1.468 vs out-of-box 1.535, rising to 1.675–1.792 with r1/r2).
7.2 Revision quality matters.
Although revised training data generally improves students, Table 3 indicates that downstream gains depend on the quality and channel-compatibility of each revision step. This is most visible for WebForm when comparing r1 and r2. Relative to org, r1 yields near-neutral WebForm changes overall (improving 4/7 student models; mean $\Delta\approx+0.01$), whereas r2 improves only 1/7 models and exhibits a negative mean change ($\Delta\approx-0.05$). The degradation is consistent across families, e.g., Llama 3.1 8B drops from $4.190$ (org) to $4.005$ (r2; $\Delta=-0.185$), Qwen3 4B from $4.010$ to $3.835$ ($\Delta=-0.175$), and Phi3.5mini 4B from $3.970$ to $3.775$ ($\Delta=-0.195$), suggesting that the additional revision step introduced in r2 is not uniformly beneficial for this channel.
Table 4 helps explain this pattern by quantifying revision success rate, where a revision is counted as successful if it fixes the targeted error or leaves the summary unchanged (thus preserving or improving quality), and as a failure if it degrades summary quality. The WebForm-specific correction used to generate r1, unn_content_webform_email_copy, is highly reliable (97% success). In contrast, the additional r2 step—a global redundancy-removal correction (unn_content_redundant) applied to all channels as the final prompt in the cascade—has substantially lower success on WebForm (61%), implying that a non-trivial fraction of WebForm training labels are worsened by this step. This failure rate is also channel-dependent: the same redundancy correction is much more reliable for TeammateChat (97%) and Email (85%), which is consistent with r2 being beneficial or neutral outside WebForm.
@@
Error type Bot Chat WebForm TeammateChat Email Overall
unn_content_requests_agent 94% - - - 94%
sentiment_inferred_frustrated 92% - - - 92%
unn_content_webform_email_copy - 97% - - 97%
unn_content_transfer - - 99% - 99%
unn_content_courtesy - - 98% 86% 92%
content_order - - - 84% 84%
unn_content_email_footer - - - 91% 91%
unn_content_redundant 78% 61% 97% 85% 80%
A second signal that revision effects are not uniform across students is Gemma3 1B: it achieves the best BotChat score with r1 ($3.520$) yet drops sharply on WebForm from $3.780$ (org) to $3.000$ (r1; $\Delta=-0.780$). Since this regression occurs prior to the r2 redundancy step, it suggests negative transfer under multi-channel fine-tuning. Many r1 edits are deletion-style and highly successful in Table 4 (e.g., 92–99% for several chat-specific errors), which can shift the target distribution toward more aggressively compressed summaries. Lower-capacity students may over-generalize this behavior across channels, harming WebForm where completeness of structured details is critical.
Overall, these results indicate that revision quality is a key determinant of ARF performance. In particular, revisions that are reliable in one channel (or on one error type) may be substantially riskier in another, and globally applied corrections such as redundancy removal should be treated as channel-dependent operations that may require tighter constraints or validation for structured inputs such as WebForms.
7.3 The better the revision, the better the fine-tuning
To quantify how revision quality translates into downstream student performance, we conducted a manual assessment of the overall summary quality change introduced by our error-correction cascade for both r1 and r2. For each intermediate output in the cascade, we compared the revised summary against the pre-prompt summary as a baseline and assigned a discrete quality-change score: a successful improvement that corrected an error was scored as +1+1, the introduction of a new error or degradation in overall quality was scored as 1-1, and revisions producing similar-quality summaries were scored as 0. We then computed the cumulative quality change for each sample by summing these scores across all prompts applied for that revision version. This procedure was performed on 100 samples per channel for r1 and r2, and Table LABEL:tab:cumulative_quality_change reports the mean cumulative quality change.
Table LABEL:tab:cumulative_quality_change shows that revision quality varies substantially across channels and between r1 and r2. BotChat exhibits a strong positive cumulative change under r1 ($0.65$), which slightly decreases under r2 ($0.59$). TeammateChat shows the largest quality gains overall, increasing from $0.58$ in r1 to $1.26$ in r2. In contrast, WebForm has a small positive cumulative change that remains unchanged between r1 and r2 ($0.12$), while Email drops sharply from $0.41$ (r1) to $0.14$ (r2).
These channel-level differences are consistent with the per-error revision reliability reported in Table 4. Redundancy removal (unn_content_redundant), the final prompt distinguishing r2 from r1, is highly reliable for TeammateChat (97% success), which aligns with the large increase in cumulative quality change for TeammateChat when moving from r1 to r2. In contrast, redundancy removal is substantially less reliable for BotChat (78%) and especially for WebForm (61%), providing a plausible explanation for why r2 yields smaller net quality improvements in BotChat and does not increase WebForm quality beyond r1. Email exhibits moderate redundancy-removal success (85%), yet the cumulative gain decreases markedly in r2, suggesting that even modest failure rates can dominate when the correction is applied broadly and interacts with other channel-specific edits.
Table 5: Cumulative summary quality change for r1/r2 from manual review
BotChat TeammateChat WebForm Email
r1 0.65 0.58 0.12 0.41
r2 0.59 1.26 0.12 0.14
Finally, we test whether this manual measure of revision quality predicts fine-tuning effectiveness. Using the cumulative quality change as a proxy for training-label quality, we performed correlation analysis against fine-tuning gains measured as $\Delta$ over org across all student models, channels, and revision versions (Figure 2). The results show a consistent positive association between revision quality and downstream improvement: Pearson correlation is $0.44$, Spearman rank correlation is $0.60$, and Kendall’s $\tau$ is $0.46$. The stronger rank-based correlations suggest that the relationship is monotonic but not strictly linear, which is plausible given model heterogeneity (capacity/architecture) and the fact that revision quality values are channel/version-specific aggregates. Collectively, these findings motivate investing in more reliable revision prompts and channel-aware validation, as gains in revision quality translate into measurable gains in student performance.
Figure 2: Correlation between training data quality and performance improvement
@@ 8 Discussion
We show that compact open-source LLMs can exceed GPT-3.5 on customer-service interaction summarization when trained with our Analyze–Revise–Finetune pipeline. In particular, Llama 3.1 8B and QWen3 4B outperform GPT-3.5 across BotChat, TeammateChat, Email, and WebForm, demonstrating that specialized production workloads can be served by smaller models with lower cost and reduced exposure of sensitive data. OpenAI does not disclose GPT-3.5’s parameter count; using GPT-3’s reported 175B parameters [2] and secondary estimates suggesting a similar scale for GPT-3.5 [16, 9], we treat our 270M–8B student models as orders of magnitude smaller than the teacher.
Our error-focused evaluation scheme ensures a consistent and systematic evaluation of summary quality, allowing for precise identification of the most common errors requiring correction. Additionally, by separating the editor LLM from the teacher LLM and utilizing Llama 3.1 70B for training data revisions, our approach further optimizes both inference cost and data privacy, offering substantial advantages compared to approaches that directly employ large teacher models for revisions.
@@ 9 Conclusion
We propose Analyze–Revise–Finetune, a targeted error-correction fine-tuning pipeline that improves training data derived from a proprietary teacher to train smaller open-source students. On a real production summarization system spanning four channels, the resulting student models match or surpass GPT-3.5 while improving cost efficiency and privacy. The ARF pipeline can be successfully extended to a wide range of LLM-based applications, offering consistent improvements in cost efficiency, output quality, and data privacy.
@@ Acknowledgments
We’d like to thank Mehran Nadjarbashi Noghani, Swapnil Thorat, Ryan Wen, Zonghan Xu, Pooja Daryani, and Megha Korade for their support in this work.
@@ Appendix A Additional Table
Unnecessary information unn_content_redundant Summary includes unnecesary information, i.e., 1) redundant or repetitive information, 2) courtesy or plesantry, 3) bot’s response, 4) customer ranting, 5) lack of specific details in the original content, 6) requets for human agent, 7) transfer to next agent, 8) email footer/signature, 9) request for email copy of original webform
@@
unn_content_platform_response_included
@@
Appendix B Additional Figures
Prompt to remove unn_content_requests_agent error from BotChat summaries Read the following summary of a chat transcript between a customer and an automatic service agent. Rewrite the summary to remove any information about the customer wanting or requesting to connect with a human agent. If there’s no other content in the summary other than customer wanting or requesting to connect with a human agent, say "<ul><li>nothing to summarize</li></ul>". Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove sentiment_inferred_frustrated error from BotChat summaries Read the following summary of a chat transcript between a customer and an automatic service agent. Rewrite the summary to remove any information about the customer’s emotion such as being frustrated, any information about the reason of the emotion, and any information about how the customer expressed their emotion. If there’s no other content in the summary other than customer’s emotion, say "<ul><li>nothing to summarize</li></ul>". Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove unn_content_courtesy error from TeammateChat summaries Read the following summary. Rewrite the summary to remove any information regarding courteous engagement between agent and customer. Courteous engagement includes expressing gratitude, assuring customer for resolution of their issue, affirming customer’s loyalty of having acocunt for many years, among others. Do not remove information regarding reassuring or confirming action if the action is conveying additional details of customer’s issue or agent’s resolution such as eBay Money Back Guarantee or timeframe. Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove unn_content_transfer error from TeammateChat summaries Read the following summary. Rewrite the summary to remove any information regarding customer being transferred to a specialist or to another agent. Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove unn_content_webform_email_copy error from WebForm summaries Read the following summary of a WebForm submitted by a customer. Rewrite the summary to remove any information about the customer wanting or requesting a copy of the communication to their email. Do not add anything else other than the revised summary.
Prompt to remove unn_content_redundant for both BotChat, TeammateChat, and WebForm summaries Read the following summary. Rewrite the summary to remove any information that is repeatedly stated. Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove content_order for Email summaries Read the following email exchanges between an eBay customer service agent and an eBay customer, along with its summary. Reorganize the summary according to the following principles: Chronological Order: Arrange events in the order they occurred—older events should appear first, followed by newer events. General to Specific: High-level descriptions of events should come first, followed by more detailed descriptions. Instructions: Rearrange the order of bullet points only. Do not rewrite or remove any bullet point. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove unn_content_courtesy for Email summaries Read the following summary. Rewrite the summary to remove any information regarding courteous engagement between agent and customer. Courteous engagement includes expressing gratitude, assuring customer for resolution of their issue, affirming customer’s royalty of having account for many years, among others. Do not remove information regarding reassuring or confirming action if the action is conveying additional details of customer’s issue or agent’s resolution such as eBay Money Back Guarantee or timeframe. Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove unn_content_email_footer for Email summaries Read the following summary. Rewrite the summary to remove any information regarding email footers or signature. Email footers or signatures includes customer’s location. Also remove any bullet point stating "User wants a copy of the communication sent to their email address.". Do not remove any important information such as Item Ids, Return Ids, Case Ids, etc or customer’s role as a buyer or a seller. Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Prompt to remove unn_content_redundant for Email summaries Read the following summary. Rewrite the summary to remove any information repeatedly stated. Do not remove any details even when there’s a general version of the same information. Instead remove general version. When the same information is stated more than twice, keep only one description. Do not remove essential details such as Item IDs, Return IDs, Case IDs, Ticket Numbers, Dates, or the customer’s role (buyer or seller). Keep the original format of the summary that is unordered list items in HTML. Do not add anything else other than the revised summary.
Figure 3: Prompts for summary error correction